You are the manager of a firm that produces Television sets and Antennas.
Information from historical sales shows that based on valuation of the two
products, there are three types of consumers, even though you are unable to
identify these consumers individually at the time of sale. In particular, you
know there are three types of consumers (1,000 of each type) with the
following valuations for the two products:
Consumer type Antennas (GHS) Television sets (GHS)
1 90 60
2 70 140
3 40 160
Suppose it costs GHS 40 to produce each television set and it costs GHS 20 to
produce each antenna.
The firm’s marketing strategists has proposed the following alternative pricing
strategies:
Strategy 1:
You charge GHS 40 for Antennas and GHS 60 for a television set.
Strategy 2
You charge GHS 90 for Antennas and GHS 160 for a television set.
Strategy 3
You charge GHS 150 for a bundle containing one antenna and one television
1
Expert's answer
2020-06-22T13:19:40-0400

Solution

Profit (∏) = R – C

Antennas (GHS)

Total Quantity = 90+70+40 = 200

Television sets (GHS)

Total Quantity = 60 + 140 + 160 = 360

Strategy 1: Charging GHS 40 for Antennas and GHS 60 for a television set.

Total Revenue (TR) = (200*40) + (360*60)

TR= 8,000 + 21,600 = 29,600

Cost (C) = 200*40 + 360*20

Cost (C) = 8,000 + 7,200 = 15,200

∏ = TR – TC = 29,600 – 15,200 = 14,400

Strategy 2: Charging GHS 90 for Antennas and GHS 160 for a television set.

Total Revenue (TR) = (200*90) + (360*160)

TR= 18,000 + 57,600 = 75,600

Cost (C) = 200*40 + 360*20

Cost (C) = 8,000 + 7,200 = 15,200

∏ = TR – TC = 75,600 – 15,200 = 60,400

Strategy 3: Charging GHS 150 for a bundle containing one antenna and one television

Total Revenue (TR) = (1*150) + (1*60)

TR= 150 + 60 = 210

Cost (C) = 1*40 + 1*20

Cost (C) = 40 + 20 = 60

∏ = TR – TC = 210 – 60 = 150